OpenStudy (anonymous):

Use the addition property and/or multiplication properties to find a and b if: -3

jhonyy9 (jhonyy9):

so if -3<x<4 and a<x-5<b so than a+5<x<b+5 ------------------------------- - so can you continue ?

OpenStudy (anonymous):

so for a: -3=a+5 so a =8 and for b: 4=b+5 so b=-1 ?

jhonyy9 (jhonyy9):

b is right but a not is check it please newly

OpenStudy (anonymous):

oh sorry, a=-8

OpenStudy (anonymous):

thanks a lot!

jhonyy9 (jhonyy9):

good luck bye